# FlashFormer: Whole-Model Kernels for Efficient Low-Batch Inference

### v1

@@ Abstract
The size and compute characteristics of modern large language models have led to an increased interest in developing specialized kernels tailored for training and inference. Existing kernels primarily optimize for compute utilization, targeting the large-batch training and inference settings. However, low-batch inference, where memory bandwidth and kernel launch overheads contribute are significant factors, remains important for many applications of interest such as in edge deployment and latency-sensitive applications. This paper describes FlashFormer, a proof-of-concept kernel for accelerating single-batch inference for transformer-based large language models. Across various model sizes and quantizations settings, we observe nontrivial speedups compared to existing state-of-the-art inference kernels. Code: https://github.com/aninrusimha/flashformer
@@ 1 Introduction
As the utility and popularity of deep learning models has increased dramatically over the past decade and a half, there has been an increasing interest in low-level kernel design, i.e. developing specialized programs tailored to optimize particular workloads on the target hardware. This has been particularly true for transformer and transformer-like architectures. This is due to a dramatic increase in task performance across domains for these models as well as the increased model scaling (i.e. increasing parameter counts and computational costs per forward pass) and inference scaling (i.e. generating more tokens) needed to achieve these results.
While most accelerator vendors develop high performance kernels for common linear algebra routines such as matrix multiplication, these general-purpose kernels are often not sufficient for LLM inference. Consequently, there has been much recent work on kernels that accelerate the inference of transformer-based LLMs, such as FlashDecoding [Dao et al., 2023] and PagedAttention [Kwon et al., 2023]. These kernels achieve efficiency gains by optimizing parts of operations common to the transformer inference pipeline. For example, FlashDecoding modifies the work partitioning in FlashAttention [Dao et al., 2022] for better parallelization during transformer decoding, while PagedAttention saves GPU memory by organizing the KV cache, thus allowing larger batch sizes.
@@
Most kernels are optimized to maximize throughput for large matrix-matrix multiplications. The key to doing so on modern accelerators is to maximize the utilization of special matrix-multiply units (MMU)—called Tensor Cores on NVIDIA hardware and MXUs on TPUs—which provide most of the computational performance of the chip. In either large-batch inference or training, this design choice optimizes the primary bottleneck of these kernels, i.e. the computational cost of matrix multiplications with large arithmetic intensity (i.e., computation per element of data movement).
However, many use cases operate under different constraints. For example, low-batch inference is primarily bottlenecked by the memory bandwidth costs of parameter and cache movement and depending on the size of the model, kernel launch overheads that arise from launching multiple separate kernels for each transformer layer. Low-batch inference workloads arise in many real-world applications, such as edge deployment, latency-sensitive applications, and more recently, inference-time scaling via longer rollouts. Moreover, low-batch inference does not obtain significant benefits from using MMUs on the hardware since the problem sizes are not bottlenecked by compute.
Writing fast kernels for these workloads requires further innovation along two common directions in kernel development research: 1) Increased operator fusion, or packing an increasing number of instructions and operations into a single kernel and 2) Increased specialization, or generating kernels that more closely optimized for the particular problem. This paper describes a strategy for writing a highly optimized inference kernel that runs the entire forward pass of a transformer-based large language model in a single kernel, and is specialized for a particular model configuration (such as the number of attention heads, embedding dimension etc) and the target hardware architecture. We use meta-programming, a shared pipelined buffer, and fast synchronization mechanisms to create a fully fused and specialized kernel. The resulting kernel, which we call FlashFormer, exploits increased overlapping of memory movement and the single fused kernel amortizes the launch overhead for the kernel over the entire forward pass. We apply this kernel to accelerate single-batch inference of Llama models [Grattafiori et al., 2024], and find that it offers respectable speedups over state-of-the-art decoding kernels.
@@ 2.1 Transformer Architecture
We target the Llama-3 [Grattafiori et al., 2024] family of models, which make use of the transformer architecture with pre-RMSNorm [Jiang et al., 2023], Grouped Query Attention [Ainslie et al., 2023], RoPE [Su et al., 2021], and SwiGLU FFN layers [Shazeer, 2020]. Each layer of a transformer model consists of six matrix operations, commonly grouped into five sublayers. Each sublayer has one matrix operation, except for the “core attention” layer, [^1] Since “attention layer” could be ambiguous, we use “core attention” to refer to the operation $\operatorname{softmax}(\alpha{\bm{Q}}{\bm{K}}^{\top}){\bm{V}}$. which has two. The matrix operations per layer with their respective input and output activations are listed in Table 1.
During autoregressive token-by-token generation, each matrix operation takes as input a large matrix stored in global memory as well as an activation computed by a previous operation. Transformer models are parameterized by their number of layers, model dimension, FFN dimension, head dimension, and the number of attention and key-value heads. The models we generate kernels for in this work are shown in Table 2.
@@
${\bm{Q}}{\bm{K}}^{\top}$ computation ${\bm{K}}[\hskip 0.85358pt:i]$ ${\bm{Q}}[i]$ ${\bm{S}}[i]$
$\bm{PV}$ computation ${\bm{V}}[\hskip 0.85358pt:i]$ $\text{softmax}(\alpha{\bm{S}}[i])$ ${\bm{O}}$
@@
Table 1: Llama matrix operations for decoding. Here $i$ is the current sequence position, and ${\bm{x}}$ is the residual stream.${\bm{W}}_{\bm{QKV}},{\bm{W}}_{{\bm{a}}_{out}}{\bm{W}}_{FFN1}$ and ${\bm{W}}_{FFN2}$ represent corresponding parameter matrices, and ${\bm{Q}},{\bm{K}}$ and ${\bm{V}}$ represent the query, key, and value matrices respectively. ${\bm{S}}$ represents the attention logits, ${\bm{P}}$ the attention weights, and ${\bm{a}}_{out}$ the output of attention. ${\bm{g}}_{inter}$ and ${\bm{g}}_{out}$ represent the output of the GLU input/gate projection and the final GLU output of the GLU layer respectively.
@@ 2.2 GPU Programming and Architecture
GPU programming optimizes over three interleaved hierarchies: the program hierarchy, the hardware hierarchy, and the memory hierarchy. The program hierarchy structures computations into kernels, thread blocks, warps, and threads to effectively expose parallelism. [^2] Note that we use the NVIDIA GPU hardware terminology in this section, other architectures use different terms for similar hardware components. The hardware hierarchy maps thread blocks onto streaming multiprocessors (SMs) which is sub-divided into multiple warp schedulers responsible for running a warp which is a collection of 32 threads [^3] This is configurable on some hardware but not on NVIDIA GPUs. executing the same instruction. Individual threads are responsible for the core computations. Finally, the memory hierarchy manages the placement of data in registers, shared memory, and global memory to minimize latency and maximize throughput. Table 3 provides an overview of the three hierarchies.
At each level except the thread level, it is possible for multiple elements of the program hierarchy to exist on the same hardware if they satisfy certain hardware and memory constraints. In particular, multiple warps can exist on one warp scheduler provided there is enough space in registers, multiple thread blocks can co-locate on an SM if there is enough shared memory and register space.
@@ 2.3 Decoding Workload and GPU Overlapping
However, this strategy does not work well for inherently sequential workloads such as the sequential computations of transformer layers. Each matrix operation in Table 1 depends on the completion of the operations before it, and each transformer layer depends on the the completion of the previous transformer layer. As a result, in normal execution there is no opportunity for overlapping between different kernels, as shown in Figure 1. Furthermore, inference kernel launch overheads can cause a nontrivial slowdown for fast decoding kernels, as shown by Spector et al. [2023].
@@ 2.4 Operator Fusion
One solution to the problem of overlapping is kernel fusion, which involves taking two or more kernels and combining them into a single kernel. Existing kernel fusions can be divided into two types. The first type fuses a single (or multiple) pointwise operations (such as an activation function) into a matrix operation. These pointwise operations are directly applied to either the input of the matrix or its output or both. These types of fusions are now commonly automated through tools such as XLA [Sabne, 2020] or PyTorch Compile [Ansel et al., 2024].
The second type involves restructuring multiple operations to remove the need for global synchronization and reduce redundant data movement between multiple matrix operations. A notable recent example of this is the FlashAttention series of kernels [Dao et al., 2022, Dao, 2024, Shah et al., 2024b], which fuses the ${\bm{Q}}{\bm{K}}^{T}$ computation, softmax calculation, and ${\bm{P}}{\bm{V}}$ computation into one kernel, reducing the memory movement from HBM to SRAM and significantly increasing arithmetic intensity over an unfused attention implementation.
@@ 3.1 General Kernel Design Principles
In this section, we elaborate on the key challenges in designing a family of whole-model kernels, and our implemented solutions.
@@
Table 4: Left: Python kernel source. Right: Generated CUDA kernel. Our metaprogramming language generates CUDA kernels from a Python interface. We use this to automate index math, perform static workload scheduling, and build kernel metaprogramming tools.
Metaprogramming for Static Optimization.
New kernels need to be implemented for each model tested. Writing such kernels manually is time intensive and prone to bugs. Implemented kernels greatly vary with both the dimensions of the models and quantization setups. To address this, we developed a metaprogramming language that, given the dimension sizes of the model, performs static workload partitioning based on predetermined rules to optimize performance. See Section 3.1 for an example. [^4] While describing this metaprogramming language in full detail is outside the scope of this paper, we believe that this language will be of interest for kernel development more generally, and intend to open source it.
Because the exact workload is known at kernel generation, we can determine at compile time what all memory operations and computations are necessary. In the memory-bound regime, we can load every element of our network parameters and KV cache exactly once. While this requires fully replicating the input, the cost of input data movement is much less than parameter and cache data movement. To efficiently generate these static workload schedules, we write a metaprogramming language in Python which automates the index math and static workload distribution and outputs CUDA kernels. As seen in Figure 2, the workload is divided among different thread blocks, then within a thread block into multiple chunks. Finally, each chunk is processed by a series of warps in a series of operations. Factors like inter-thread block workload partitioning, chunk sizes, and intra chunk workload partitioning are optimized to maximize utilization of hardware resources, especially global-to-shared memory bandwidth.
Memory pipeline.
Each layer has different sized matrices and different shared memory requirements. As shown in Figure 2, we design and implement a unified memory pipeline that allows for the memory movement of arbitrary layers to overlap, and for memory movement to overlap all other operations. This overlapping of memory movement with computation is also referred to as a producer-consumer pipeline design and is used by generally used by fast GEMM-based kernels like FlashAttention-3 for Hopper GPUs [Shah et al., 2024a].
Specifically, each matrix listed in Table 1 is split into chunks of a fixed size in bytes. This chunk size remains the same throughout the model, although the data type of the chunk, number of elements, and matrix dimensions are subject to change. Each matrix’s chunks are split among thread blocks at compile time. While all weight matrix chunks are loaded every iteration, KV cache chunks are only loaded as needed. Due to this structure, different sections of the matrix buffer can store matrix chunks from different layers or sublayers. The number of matrix buffers stored at a given time is the “depth” of the pipeline, and the size of one matrix buffer is the size of each pipeline stage.
Global memory ordering.
One problem with our approach is that the inputs for subsequent layers rely on the outputs of prior layer. Given the relaxed memory ordering of CUDA computations, we use a per thread global memory fence coupled with an atomic-based global synchronization mechanism to synchronize volatile reads and writes. This application of a traditional atomic synchronization ensures all participating threads’ global memory operations are visible to all other participating threads before continuing execution to avoid a read-after-write race condition. While atomic-based synchronization is far from novel, we believe we are the first to use it to combine different transformer layers into a single fused kernel.
3.2 Transformer-specific Kernel Design
Figure 3: Conceptual diagram of a standard attention forward pass and our method. Left: a standard attention forward pass in an inference pipeline, broken down into four kernels. Right: FlashFormer, where all operations are part of one kernel invocation. A detailed diagram of our attention setup can be found in Figure 5.
Attention.
We divide the attention sublayer into three parts, each of which has synchronous global communication between them: QKV Projection, Core Attention, and Attention Output Projection. A comparison of standard workload partitioning to ours can be found in Figure 3. A detailed diagram of the operations within each of these parts can be seen in Figure 5 of the appendix.
Both QKV Projection and Attention Output projection are matrix vector operations in the batch one case. For these operations, we cache the input in shared memory, and each chunk of the matrix corresponds to some number of rows of the weight matrix. We compute RMSNorm, RoPE, and then QKV projection, and add the outputs of Attention Output projection directly to the residual stream.
For Core Attention, we use a modified version of the FlashDecoding kernel [Dao et al., 2023]. Thread blocks are split among the different KV heads. For each KV head, each SM retains a replicated copy of queries which multiply with that particular KV head. The KV cache is then partitioned into matrix chunks, which are fed into the memory buffers. One problem with our pipelined approach is that we have no memory ordering that guarantees the elements of the KV cache correspond to the current token. We thus synchronously load the keys and values for the current token in.
After the entire KV cache has been read and the partial core attention output computed, we perform a reduction per KV head before writing to a shared buffer. The reduction occurs in three stages. First, each thread block writes required metadata for the FlashDecoding reduction, e.g. the maximum logit and the softmax denominator. In the second stage, each thread block reads the combined metadata and computes the global softmax denominator. It then rescales its output by the ratio of the thread block’s softmax denominator to the combined softmax denominator, as in Milakov and Gimelshein [2018]. Finally, all threads add their rescaled outputs to a cumulative sum in global memory. As opposed to the normal FlashDecoding kernel, the reduction is fused into the output computation kernel, removing one activation round trip.
GLU.
In our kernel, the GLU in/gate operation is a matrix vector operation very similar to the QKV and Attention Output Projection operations in Attention, as shown in Figure 6. Unlike those operators, the output of each chunk of the GLU in/gate operation is not written to global memory.
Instead, we compute the SwiGLU activation on the outputs of the GLU in/gate function and then multiply it with a matrix chunk of GLU out. We then keep a vector of size $d_{model}$ for accumulating partial outputs. With this approach, we can iterate through the entirety of the GLU computation without needing to do any synchronization or reduction. At the end the computation, we reduce all of these partial outputs along with the input.
4 Experiments & Results

### v2

@@ Abstract
The size and compute characteristics of modern large language models have led to an increased interest in developing specialized kernels tailored for particular training and inference workloads. Existing kernels primarily optimize for compute utilization, targeting the large-batch training and inference settings. However, low-batch inference, where memory bandwidth and kernel launch overheads are significant factors, remains important for many applications of interest such as in edge deployment and latency-sensitive applications. This paper describes FlashFormer, which fuses the entire transformer forward pass into a single kernel for accelerating low-batch inference of large language models. Across various model sizes and quantizations settings, FlashFormer achieves nontrivial speedups compared to existing inference kernels.
footnotetext: 1Massachusetts Institute of Technology, Cambridge, MA, USA 2MIT-IBM Watson AI Lab, Cambridge, MA, USA. Correspondence to: Aniruddha Nrusimha <anin@mit.edu>.
Preprint.
@@ 1 Introduction
As the utility and popularity of deep learning models has steadily increased, there has been an increasing interest in low-level kernel design, i.e., developing specialized programs tailored to optimize particular workloads on the target hardware. This has been particularly true for large language models (LLMs) based on the transformer architecture, due to a dramatic increase in task performance across domains for these models obtained from model scaling (i.e., increasing parameter counts and computational costs per forward pass) and inference scaling (i.e., generating more tokens in parallel or serially).
While most accelerator vendors develop high performance kernels for common linear algebra routines such as matrix multiplications, these general-purpose kernels are often not sufficient for LLM inference. Consequently, there has been much recent work on kernels that accelerate the inference of transformer-based LLMs, such as FlashDecoding (Dao et al., 2023) and PagedAttention (Kwon et al., 2023). These kernels achieve efficiency gains by optimizing parts of operations common to the transformer inference pipeline. For example, FlashDecoding modifies the work partitioning in FlashAttention (Dao et al., 2022; Dao, 2024) for better parallelization during transformer decoding, while PagedAttention saves GPU memory by organizing the KV cache, thus allowing larger batch sizes.
@@
Most existing kernels are optimized to maximize throughput for large matrix-matrix multiplications (matmuls). The key to doing so on modern accelerators is to maximize the utilization of special matrix-multiply units (MMU)—called Tensor Cores on NVIDIA hardware and MXUs on TPUs—which provide most of the computational performance of the chip. In either large-batch inference or training, this design choice optimizes the primary bottleneck of these kernels, i.e., the computational cost of matmuls with high arithmetic intensity (computation per element of data movement).
However, many use cases operate under different constraints. For example, low-batch inference is primarily bottlenecked by the memory bandwidth costs of parameter and KV-cache movement; depending on the size of the model, kernel launch overheads that arise from launching multiple separate kernels for each transformer layer can also be a bottleneck. Low-batch inference workloads arise in many real-world applications, such as edge deployment, latency-sensitive applications, and more recently, inference-time scaling via longer rollouts. Moreover, low-batch inference does not obtain significant benefits from using MMUs on the hardware since the problem sizes are not bottlenecked by compute.
Writing fast kernels for these workloads requires further innovation along two common directions in kernel development research: 1) Increased operator fusion, or packing an increasing number of instructions and operations into a single kernel and 2) Increased specialization, or generating kernels that more closely optimized for the particular problem. This paper describes a strategy for writing a highly optimized inference kernel that runs the entire forward pass of a transformer-based LLM in a single kernel, and is specialized for a particular model configuration (such as the number of attention heads, embedding dimension, etc) and the target hardware architecture. We use meta-programming, a shared pipelined buffer, and fast synchronization mechanisms to create a fully fused and specialized kernel. The resulting kernel, which we call FlashFormer, exploits increased overlapping of memory movement and the single fused kernel amortizes the launch overhead for the kernel over the entire forward pass. We apply this kernel to accelerate low-batch inference of Llama models (Grattafiori et al., 2024), and find that it offers consistent speedups over existing decoding kernels.
@@ 2.1 Transformer Architecture
We target the Llama (Grattafiori et al., 2024) family of transformer architectures, which make use of the pre-RMSNorm (Jiang et al., 2023), Grouped Query Attention (Ainslie et al., 2023), RoPE (Su et al., 2021), and SwiGLU FFN layers (Shazeer, 2020). Each layer of a transformer model consists of six matrix operations, commonly grouped into five sublayers. Each sublayer has one matrix operation, except for the “core attention” layer, [^1] Since “attention layer” could be ambiguous, we use “core attention” to refer to the operation $\operatorname{softmax}(\alpha{\bm{Q}}{\bm{K}}^{\top}){\bm{V}}$. which has two. The matrix operations per layer with their respective input and output activations are listed in Table˜1. These components are common to many other recent families of LLMs.
During autoregressive token-by-token generation, each matrix operation takes as input a large matrix stored in global memory as well as an activation computed by a previous operation.
Transformer LLMs are parameterized by the number of layers, model dimension, FFN dimension, head dimension, and the number of attention and key-value heads. The models we generate kernels for are shown in Table˜2.
@@
${\bm{Q}}{\bm{K}}^{\top}$ computation ${\bm{K}}[\hskip 0.85358pt:\hskip-1.99168pti]$ ${\bm{Q}}[i]$ ${\bm{S}}[i]$
$\bm{PV}$ computation ${\bm{V}}[\hskip 0.85358pt:\hskip-1.99168pti]$ $\text{softmax}(\alpha{\bm{S}}[i])$ $\bm{O}$}
@@
[…]
@@ 2.2 GPU Programming and Architecture
GPU programming optimizes over three interleaved hierarchies: the program hierarchy, the hardware hierarchy, and the memory hierarchy. The program hierarchy structures computations into kernels, thread blocks, warps, and threads to effectively expose parallelism. [^2] We use the NVIDIA GPU hardware terminology throughout this work; other architectures use different terms for similar hardware components. The hardware hierarchy maps thread blocks onto streaming multiprocessors (SMs) which is sub-divided into multiple warp schedulers responsible for running a warp which is a collection of 32 threads executing the same instruction. Individual threads are responsible for the core computations. Finally, the memory hierarchy manages the placement of data in registers, shared memory, and global memory to minimize latency and maximize throughput. Table˜3 provides an overview of the three hierarchies.
At each level except the thread level, it is possible for multiple elements of the program hierarchy to exist on the same hardware if they satisfy certain hardware and memory constraints. In particular, multiple warps can exist on one warp scheduler provided there is enough space in registers, andmultiple thread blocks can co-locate on an SM if there is enough shared memory and register space.
@@ 2.3 Decoding Workload and GPU Overlapping
However, this strategy does not work well for inherently sequential workloads such as the sequential computations of transformer layers. Each matrix operation in Table˜1 depends on the completion of the operations before it, and each transformer layer depends on the the completion of the previous transformer layer. As a result, in normal execution there is no opportunity for overlapping between different kernels, as shown in Figure˜1. Furthermore, inference kernel launch overheads can cause a nontrivial slowdown for fast decoding kernels, as noted by Spector et al. (2023).
@@ 2.4 Operator Fusion
One solution to the problem of overlapping is kernel fusion, which involves taking two or more kernels and combining them into a single kernel. Existing kernel fusions can be divided into two types. The first type fuses a single (or multiple) pointwise operations (such as an activation function) into a matrix operation. These pointwise operations are directly applied to either the input of the matrix or its output or both. These types of fusions are now commonly automated through tools such as XLA (Sabne, 2020) or PyTorch Compile (Ansel et al., 2024).
The second type involves restructuring multiple operations to remove the need for global synchronization and reduce redundant data movement between multiple matrix operations. A notable recent example of this is the FlashAttention series of kernels (Dao et al., 2022; Dao, 2024; Shah et al., 2024b), which fuses the ${\bm{Q}}{\bm{K}}^{T}$ computation, softmax calculation, and ${\bm{P}}{\bm{V}}$ computation into one kernel, reducing the memory movement from HBM to SRAM and significantly increasing arithmetic intensity over an unfused kernel.
3 Preliminary Study with Deep Linear Networks
We motivate the development of whole-model kernels for transformer inference by demonstrating the latency benefits in the case of deep linear networks. Insofar as deep linear networks make use of simpler primitives (i.e., just a stack of linear layers), speedups would be difficult to realize; thus, any speedups observed in this setting would provide promising motivation for fusing the entire transformer forward pass. We measure the latency reduction from eliminating kernel overheads at a variety of widths and depths, and ablate the latency reduction from overlapping memory movement across layers. We measure both the execution time and achieved memory bandwidth; since our workloads are memory-bound, higher memory bandwidth indicates better hardware utilization.
Kernel fusion across layers.
We first study fusing the linear layers into a single kernel thus eliminating kernel overheads. We use a cooperative square linear kernel with a dynamic number of layers with a batch-one input (i.e., a vector). As a cooperative kernel, only one thread block runs per SM. Each thread block works on a predetermined portion of the workload, and performs a global sync in between layers. For different problem sizes (model dimension / number of layers), we perform autotuning across the number of consumer warps, number of pipeline stages, and size of each stage of the memory pipeline. The latency is shown in Table˜4, and the memory bandwidth is shown in Figure˜2. We find that kernel fusion across layers improves realized memory bandwidth. This effect is larger for smaller layers and diminishes as the number of fused layers increases.
Dim Layers Baseline Fused + Overlap
1 0.013 0.013 0.012
2K 4 0.050 0.029 0.027
32 0.403 0.169 0.154
1 0.021 0.021 0.021
4K 4 0.084 0.060 0.060
32 0.675 0.423 0.423
1 0.062 0.062 0.062
8K 4 0.250 0.221 0.218
32 1.998 1.543 1.513
Table 4: Latencies of stacked linear layers, measured in ms. Baseline refers to stacking separate kernel launches, Fused combines the forward pass into a single kernel. + Overlap is same as fused but makes use of cross layer overlapping of memory movement.
Memory overlapping.
A benefit of working with a single kernel is that we can initiate memory movement for future layers before the computation for current layers finishes. The additional benefits of memory overlapping are shown in Table˜4, representing an additional few percentage points of latency improvement over the fused kernel.
Figure 2: Memory bandwidth achieved by fused stacked linear layers. Cross layer overlapping refers to starting to load weights for future layers before computation for the current layer has finished.
These promising initial results motivate the development of whole-model kernels for transformer inference.
@@ 4.1 General Kernel Design Principles
We first elaborate on the key challenges in designing a family of whole-model kernels and describe how FlashFormer overcomes these challenges.
Metaprogramming for static optimization.
New kernels need to be implemented for each model tested. Writing such kernels manually is time intensive and prone to bugs. Implemented kernels can vary greatly depending on factors such as model dimension and quantization settings. To address this, we developed a metaprogramming language (in Python) called Cheetah [^3] https://github.com/cheetah-lang, which, given the dimension sizes of the model, performs static workload partitioning based on predetermined rules.
Figure 4: Conceptual diagram of a standard attention forward pass and our method. Left: a standard attention forward pass in an inference pipeline, broken down into four kernels. Right: FlashFormer, where all operations are part of one kernel invocation.
Figure 5: The Attention forward pass per layer. Large matrices (green) are loaded asynchronously from global, and do not wait on warps performing computation. The consumers works synchronously, with 3 global synchronizations per attention forward pass. Each of the repeated computations (in red) corresponds to a asynchronously loaded matrix split into chunks.
Concretely, because the exact workload is known at kernel generation, we can determine at compile time what memory operations and computations are necessary. In the memory-bound regime, we can load every element of our network parameters and KV cache exactly once. While this requires fully replicating the input, the cost of input data movement is orders of magnintude less than parameter and cache data movement. Cheetah automates the index math and static workload distribution and outputs CUDA kernels, thus efficiently generating these static workload schedules. An example of the metaprogamming language, along with a generated kernel, is provided in Section˜4.2, with more advanced examples in LABEL:tab:cheetah_advanced in the appendix.
As seen in Figure˜3, the workload is divided among different thread blocks, then within a thread block into multiple chunks. Finally, each chunk is processed by a series of warps in a series of operations. Factors like inter-thread block workload partitioning, chunk sizes, and intra chunk workload partitioning are optimized to maximize utilization of hardware resources, especially global-to-shared memory bandwidth.
Memory pipeline.
Each layer has different sized matrices and different shared memory requirements. As shown in Figure˜3, we design and implement a unified memory pipeline that allows for the memory movement of arbitrary layers to overlap, and for memory movement to overlap all other operations. This overlapping of memory movement with computation is also referred to as a producer-consumer pipeline design and is used by generally used by fast GEMM-based kernels like FlashAttention-3 for Hopper GPUs (Shah et al., 2024a).
Specifically, each matrix listed in Table˜1 is split into chunks of a fixed size in bytes. This chunk size remains the same throughout the model, although the data type of the chunk, number of elements, and matrix dimensions are subject to change. Each matrix’s chunks are split among thread blocks at compile time. While all weight matrix chunks are loaded every iteration, KV cache chunks are only loaded as needed. Due to this structure, different sections of the matrix buffer can store matrix chunks from different layers or sublayers. The number of matrix buffers stored at a given time is the “depth” of the pipeline, and the size of one matrix buffer is the size of each pipeline stage.
Global memory ordering.
One potential issue with our approach is that the inputs for subsequent layers rely on the outputs of prior layer. Given the relaxed memory ordering of CUDA computations, we use a per thread global memory fence coupled with an atomic-based global synchronization mechanism to synchronize volatile reads and writes. This application of a traditional atomic synchronization ensures all participating threads’ global memory operations are visible to all other participating threads before continuing execution to avoid a read-after-write race condition. While atomic-based synchronization is far from novel, using it to combine different transformer layers into a single fused kernel has not been explored before to our knowledge.
4.2 Transformer-specific Kernel Design
Attention.
We divide the attention sublayer into three parts, each of which has synchronous global communication between them: QKV Projection, Core Attention, and Attention Output Projection. A comparison of standard workload partitioning to ours can be found in Figure˜4. Figure˜5 delineates a more detailed diagram of the operations within each of the above parts.
In the batch one case, both QKV Projection and Attention Output projection are matrix vector operations. For these operations, we cache the input in shared memory, and each chunk of the matrix corresponds to some number of rows of the weight matrix. We compute RMSNorm, RoPE, and then QKV projection, and add the outputs of Attention Output projection directly to the residual stream.
For Core Attention, the thread blocks are split among the different KV heads, and for each KV head, each SM retains a replicated copy of queries which multiply with that particular KV head. The KV cache is then partitioned into matrix chunks, which are fed into the memory buffers. Since our pipelined approach has no memory ordering that guarantees the elements of the KV cache correspond to the current token, we synchronously load the keys and values for the current token.
After the entire KV cache has been read and the partial core attention output computed, we perform a reduction per KV head before writing to a shared buffer. The reduction occurs in three stages. First, each thread block writes required metadata for the FlashDecoding-like reduction, e.g. the maximum logit and the softmax denominator. In the second stage, each thread block reads the combined metadata and computes the global softmax denominator. It then rescales its output by the ratio of the thread block’s softmax denominator to the combined softmax denominator, as in Milakov & Gimelshein (2018). Finally, all threads add their rescaled outputs to a cumulative sum in global memory. Compared to the FlashDecoding kernel (Dao et al., 2023), the reduction is fused into the output computation kernel, removing one activation round trip.
GLU.
In our kernel, the GLU in/gate operation is a matrix vector operation very similar to the QKV and Attention Output Projection operations in Attention, as shown in Figure˜6. Unlike those operators, the output of each chunk of the GLU in/gate operation is not written to global memory. Instead, we compute the SwiGLU activation on the outputs of the GLU in/gate function and then multiply it with a matrix chunk of GLU out. We then keep a vector of size $d_{model}$ for accumulating partial outputs. With this approach, we can iterate through the entirety of the GLU computation without needing to do any synchronization or reduction. At the end the computation, we reduce all of these partial outputs along with the input. See Figure˜6 for a full schematic.
@@ 4.2 Transformer-specific Kernel Design
Table 5: Upper Left: Python kernel source. Upper Right: Generated CUDA kernel. Our metaprogramming language generates CUDA kernels from a Python interface. We use this to automate index math, perform static workload scheduling, and build kernel metaprogramming tools. More advanced examples can be found in Table˜7 in the appendix.
Model Prefill Tokens per second
Size Size GPTFast vLLM FlashFormer
8B 128 168 146 184
(BF16) 1024 159 144 182
3072 158 141 180
6144 143 138 176
8B 128 290 193 311
(INT4) 1024 284 186 310
3072 266 189 307
6144 228 175 301
70B 128 47 - 48
(INT4) 1024 45 - 48
3072 43 - 47
6144 40 - 45
Table 6: Prefill size and tokens per second (TPS) for the model configurations in the batch one case for the LLama 3.1 series of models. We were unable to get vLLM numbers for LLama 3.1 70B quantized results due to system errors.